OpenStudy (anonymous):

Does someone know a shortcut for finding the coffiecient of the product terms in the binomial expansion?? eg: Cofficient of x^5 in the expansion of ((1+x)^6 (3-4x)^8) ??

ganeshie8 (ganeshie8):

im not sure of any shortcuts, but u could simplify the problem a little bit... you only need to consider : (1+x)^5(3-4x)^5

ganeshie8 (ganeshie8):

5 = 0+5 = 1+4 = 2+3 = 3+2 = 4 + 1 = 5+ 0

ganeshie8 (ganeshie8):

so that gives 6 additions which is really not what you're looking for

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

This is the general way ...anyways thank you!!

ganeshie8 (ganeshie8):

you familiar wid below thing right ? \[^nC_0 + ^n C_1 + ^nC_2 +... + ^nC_n = 2^n\]

ganeshie8 (ganeshie8):

we may use it and derive some formula i think

OpenStudy (anonymous):

kk